OpenStudy (anonymous):

Two trains leave the station at the same time, one heading west and the other east. The westbound train travels at 90 kilometers per hour. The eastbound train travels at 80 kilometers per hour. How long will it take for the two trains to be 442 kilometers apart? Do not do any rounding.

OpenStudy (welshfella):

the speed at which they are going apart is 90 + 80 = 170 mph to find the time use the formula speed = distance / time rearranging time = distance / speed

OpenStudy (welshfella):

just plug in speed and distance

OpenStudy (welshfella):

ok? any questions?

OpenStudy (welshfella):

time = 442 / 170 = ?

OpenStudy (anonymous):

2.6 hours

OpenStudy (anonymous):

thanks @welshfella